OpenStudy (anonymous):

When using the Babylonian method to approximate √5 if you start with a guess of 2, then what will be your next guess? A. 2.25 B. 2.5 C. 3 D. 2

OpenStudy (blockcolder):

The Babylonian method is \[x_{n+1}=\frac{1}{2}\left(x_n+\frac{a}{x_n}\right)\] where a is the number whose square root you're looking for, right?

OpenStudy (blockcolder):

Since a=5 and x_0=2, your next guess is \[x_1=\frac{1}{2}\left(2+\frac{5}{2}\right)\] If you plug this in your calculator, you should get the right answer.